OpenStudy (katie_rogers101):

Hi this question really confused me. I tried to answer it but I got it wrong can anyone help? What is the side length, if the area of the square is: 9. 64x sqrt(y) 8x^2 fourthroot y

OpenStudy (katie_rogers101):

That was my wrong answer

OpenStudy (mathstudent55):

Is 9 part of the area?

OpenStudy (mathstudent55):

If 9 is not part of the problem, then you have: \(A = 64x\sqrt y\) \(s^2 = 64x \sqrt y\) \(s = \sqrt{64x \sqrt y} \) \(s = \sqrt {64} \sqrt x \sqrt{\sqrt y} \) \(s = 8 \sqrt x \sqrt[4] y\) It looks like your only mistake was that you squared x instead of taking its square root.

OpenStudy (katie_rogers101):

oh okay well thank you. I thought I was messing up on the 4throot thing